If the sum of a number and five is doubled, the result is three less than the number.. find the number
Found 2 solutions by macston, fractalier:
Answer by macston(5194)

.
x= the number
.
2(x+5)=x-3
2x+10=x-3
x+10=-3
x=-13
.

Answer by fractalier(6550)

Call the number x.
Then we can write
(x+5)*2 = x - 3
2x + 10 = x - 3
Now subtract x and subtract 10 and get
x + 10 = -3
x = -13